# Feature Re-Embedding: Towards Foundation Model-Level Performance in Computational Pathology

### v3

@@ Abstract
Multiple instance learning (MIL) is the most widely used framework in computational pathology, encompassing sub-typing, diagnosis, prognosis, and more. However, the existing MIL paradigm typically requires an offline instance feature extractor, such as a pre-trained ResNet or a foundation model. This approach lacks the capability for feature fine-tuning within the specific downstream tasks, limiting its adaptability and performance. To address this issue, we propose a Re-embedded Regional Transformer (RT) for re-embedding the instance features online, which captures fine-grained local features and establishes connections across different regions. Unlike existing works that focus on pre-training powerful feature extractor or designing sophisticated instance aggregator, RT is tailored to re-embed instance features online. It serves as a portable module that can seamlessly integrate into mainstream MIL models. Extensive experimental results on common computational pathology tasks validate that: 1) feature re-embedding improves the performance of MIL models based on ResNet-50 features to the level of foundation model features, and further enhances the performance of foundation model features; 2) the RT can introduce more significant performance improvements to various MIL models; 3) RT-MIL, as an RT-enhanced AB-MIL, outperforms other latest methods by a large margin. The code is available at: https://github.com/DearCaat/RRT-MIL.
@@ 1 Introduction
In this paper, we propose Re-embedded Regional Transformer (RT), which leverages the advantages of the native MSA while overcoming its limitations. Specifically, RT applies the native MSA to each local region separately. Then, it uses a Cross-region MSA (CR-MSA) to fuse the information from different regions. Finally, a novel Embedded Position Encoding Generator (EPEG) is used to effectively encode the positional information of the patches. By incorporating with mainstream MIL models, the proposed RT can ensure efficient computation while maintaining powerful representation capabilities to fine-tune the offline features according to the specific downstream tasks. The main contributions can be summarized as follows:
@@
For re-embedding instance features, we design a Re-embedded Regional Transformer (RT) which can be seamlessly integrated into mainstream MIL models to further improve performance. By incorporating the RT into AB-MIL, we present the RT-MIL, which achieves state-of-the-art performance on various computational pathology benchmarks.
We introduce two novel components for the RT: the Cross-region MSA and the Embedded Position Encoding Generator. The former enables effective information fusion across different regions. The latter combines the benefits of relative and convolutional position encodings to encode the positional information more effectively.
@@ 3.2 Re-embedded Regional Transformer
As illustrated in Figure 2, we propose a Re-embedded Regional Transformer (RTransformer, RT) to re-embed the input instance features as the new instance representations,
where is the mapping function of the RTransformer here, is the processed input instance features, and is the new embedding of the -th instance. is the dimension of the embedding. The RTransformer can be flexibly plugged into the MIL framework as a re-embedding module after feature input and before instance aggregation to reduce bias caused by the shift between offline feature learning and downstream tasks. The whole re-embedding process in the RTransformer can be formulated as,
@@
| Methods | CAMELYON-16 | TCGA-BRCA | |||||
| Accuracy | AUC | F1-score | Accuracy | AUC | F1-score | ||
| ResNet-50 ImageNet pretrained | AB-MIL [14] | 90.060.72 | 94.540.30 | 87.830.83 | 86.414.92 | 91.102.52 | 81.644.71 |
| CLAM [22] | 90.140.85 | 94.700.76 | 88.100.63 | 85.172.70 | 91.671.78 | 80.373.04 | |
| DSMIL [18] | 90.171.02 | 94.570.40 | 87.651.18 | 87.202.69 | 91.581.33 | 82.412.92 | |
| TransMIL [26] | 89.222.32 | 93.512.13 | 85.104.33 | 84.682.67 | 90.801.91 | 79.862.63 | |
| DTFD-MIL [45] | 90.220.36 | 95.150.14 | 87.620.59 | 85.921.76 | 91.431.64 | 81.092.05 | |
| IBMIL [20] | 91.230.41 | 94.801.03 | 88.800.89 | 84.193.40 | 91.012.32 | 79.453.42 | |
| MHIM-MIL [30] | 91.810.82 | 96.140.52 | 89.940.70 | 86.735.59 | 92.361.58 | 82.435.47 | |
| RT-MIL | 92.400.31 | 97.320.29 | 90.630.45 | 88.330.67 | 93.171.45 | 83.700.95 | |
| PLIP WSI pretrained | AB-MIL [14] | 94.660.42 | 97.300.31 | 93.290.54 | 85.452.32 | 91.732.26 | 80.602.66 |
| CLAM [22] | 93.730.54 | 97.170.50 | 91.600.60 | 86.701.35 | 92.162.02 | 81.911.78 | |
| DSMIL [18] | 94.400.85 | 97.060.56 | 92.781.15 | 87.252.70 | 91.801.67 | 82.182.28 | |
| TransMIL [26] | 94.400.43 | 97.880.21 | 92.810.43 | 85.833.44 | 92.172.20 | 81.123.25 | |
| DTFD-MIL [45] | 94.570.31 | 97.290.23 | 93.120.40 | 86.422.67 | 92.162.42 | 81.772.73 | |
| IBMIL [20] | 93.900.66 | 97.040.18 | 92.440.64 | 87.571.48 | 91.711.74 | 82.782.02 | |
| MHIM-MIL [30] | 95.320.31 | 97.790.15 | 94.130.42 | 87.072.20 | 93.172.00 | 82.482.50 | |
| RT-MIL | 95.490.00 | 98.050.29 | 94.290.04 | 88.823.22 | 93.801.24 | 84.553.55 | |
@@ 3.3 RTransformer-based MIL
where denotes any MIL loss. RTransformer-based MIL (RT-MIL) adopts the instance aggregation method and the bag classifier of AB-MIL [14] by default.
@@ 4.2.1 Cancer Diagnosis, and Sub-typing
Table 1 presents the diagnosis and sub-typing performances of different MIL approaches on the C16 and BRCA datasets. The results demonstrate that our proposed RT-MIL achieves the best performance under all metrics on all benchmarks. Specifically, RT-MIL gets 0.59%, 1.18%, and 0.69% performance gains over the second-best methods in Accuracy, AUC, and F1-score respectively on the C16 dataset. On the BRCA dataset, the AUC improvement is 0.69%. RT-MIL employs the same aggregation and classification methods as AB-MIL. However, RT-MIL significantly outperforms AB-MIL. It increases the AUC by 2.78% and 1.77% on the C16 and BRCA datasets, respectively. The sub-typing results on NSCLC in Table 2 support a similar observation. We attribute these substantial performance improvements to the additional re-embedding step based on our proposed RT, which surpasses the performance of the foundation model (+0.02% AUC on C16, +1.37% on BRCA, +0.72% on NSCLC). In addition, we find that RT can further enhance the features of the foundation model, achieving considerable improvement. This validates the effectiveness of re-embedding.
@@
| ResNet-50 | AB-MIL | 90.32 | 95.29 | 89.83 |
| CLAM | 90.52 | 95.37 | 90.08 | |
| DSMIL | 90.43 | 95.60 | 90.03 | |
| TransMIL | 90.04 | 94.97 | 89.94 | |
| DTFD-MIL | 89.85 | 95.55 | 89.60 | |
| IBMIL | 90.04 | 95.57 | 89.73 | |
| MHIM-MIL | 91.27 | 96.02 | 90.85 | |
| RT-MIL | 91.75 | 96.40 | 91.26 | |
| PLIP | AB-MIL | 90.99 | 95.68 | 90.52 |
| CLAM | 90.80 | 95.46 | 90.38 | |
| DSMIL | 90.60 | 95.78 | 90.24 | |
| TransMIL | 89.09 | 95.30 | 88.83 | |
| DTFD-MIL | 90.42 | 95.83 | 89.91 | |
| IBMIL | 91.18 | 95.62 | 90.94 | |
| MHIM-MIL | 91.74 | 96.21 | 91.20 | |
| RT-MIL | 92.13 | 96.40 | 91.83 | |
@@
| ResNet-50 | AB-MIL | 57.50 | 58.78 | 56.51 |
| CLAM | 57.57 | 59.60 | 56.65 | |
| DSMIL | 57.42 | 59.31 | 55.03 | |
| TransMIL | 58.90 | 64.11 | 56.39 | |
| DTFD-MIL | 56.98 | 59.48 | 55.16 | |
| IBMIL | 58.41 | 58.58 | 59.18 | |
| MHIM-MIL | 58.36 | 60.32 | 56.08 | |
| RT-MIL | 61.13 | 67.19 | 60.95 | |
| PLIP | AB-MIL | 59.18 | 62.09 | 57.12 |
| CLAM | 61.58 | 64.05 | 58.00 | |
| DSMIL | 58.96 | 63.82 | 56.12 | |
| TransMIL | 56.20 | 63.55 | 58.84 | |
| DTFD-MIL | 59.67 | 61.78 | 57.75 | |
| IBMIL | 56.32 | 58.86 | 57.33 | |
| MHIM-MIL | 60.92 | 62.94 | 55.95 | |
| RT-MIL | 63.98 | 65.94 | 60.42 | |
@@ 4.2.2 Survival Prediction
Table 3 shows the experimental results on three survival prediction datasets. It is worth noting that our proposed RT-MIL model demonstrates outstanding performance, attaining a C-index of 61.13% on the BLCA, 67.19% on the LUAD, and 60.95% on the LUSC. It outperforms the compared methods by a significant margin, with improvements of 2.23%, 3.08%, and 1.77% over the second-best methods, respectively. Furthermore, our proposed feature re-embedding strategy can yield substantial improvements even when working with high-quality features extracted by the foundation model. Particularly, compared to AB-MIL, our feature re-embedding strategy brings performance improvements of 4.8%, 3.85%, and 3.3% for the three datasets, respectively. These results highlight the consistent and reliable performance of our proposed strategy and method, indicating its efficacy in predicting survival outcomes.
@@ 4.3.1 Re-Embedding Matters in Pathology
Foundation Model Features vs. Re-embedding Features: Table 4 and Figure 4 compare the performance of the features extracted by the foundation model PLIP [13] and various re-embedding features. The PLIP is based on the multi-modal foundation model CLIP [24] and uses up to 200K slide-text pairs for pre-training. Although this high-cost pre-training brings some improvement on different tasks, it still has bottlenecks. We attribute this to the dilemma of the traditional paradigm that even the best offline pre-training features cannot address the issue of insufficient feature fine-tuning for downstream tasks. In contrast, re-embedding modules that can be end-to-end trained with MIL models provide supervised feature fine-tuning, which enables full exploitation of the knowledge beneficial to the final task. Hence, we can see from Table 4 that any re-embedding structure achieves a considerable improvement on all tasks. Some tailored structures, such as our proposed RT, can have a significant performance advantage in most of the tasks. Moreover, Figure 4 shows that this improvement is not limited to the classical AB-MIL, but can widely benefit different MIL models. Table 1 demonstrates that the re-embedding approach is still effective on foundation model features. Therefore, compared to foundation model features, re-embedding features are a cheaper, more versatile alternative and effective booster.
@@
| +RT (local) | 97.32 | 96.40 | 67.19 | 6.5s |
Different Re-embedding Discussion: The bottom part of Table 4 presents the performance of AB-MIL under different feature settings. We employ four methods, including TranMIL, N-MSA, the local version of N-MSA, and our proposed RT to re-embed the features. From observations, all four re-embedded AB-MILs perform better than the original one. The performance improvement by TransMIL, N-MSA, and RT are 1.28%, 1.68%, and 2.80%, respectively, on C16, while these numbers on LUAD are 4.46%, 4.14%, and 8.41%, respectively. This phenomenon validates the importance of re-embedding in MIL-based computational pathology. Among the four employed re-embedding approaches, RT boosts the AB-MIL with the most considerable improvements while incurring the lowest computational cost. Specifically, RT achieves more performance gains (+0.85% on C16 and +2.63% on LUAD) while only requiring a 1/5 inference time compared with the second-best approaches.
The Applicability of RT in MIL Frameworks: We incorporate the RT into different MIL frameworks as a re-embedding module for studying its applicability. The performance improvements achieved by this re-embedding module in different frameworks are shown in Figure 4. The results reveal that the RT is capable of improving all MIL frameworks on all tasks. Moreover, the improvement brought by RT surpasses the foundation model features on two tasks except for diagnosis, and reaches a similar level on C16. This clearly verifies the good applicability of RT.

### v4

@@ Abstract
Multiple instance learning (MIL) is the most widely used framework in computational pathology, encompassing sub-typing, diagnosis, prognosis, and more. However, the existing MIL paradigm typically requires an offline instance feature extractor, such as a pre-trained ResNet or a foundation model. This approach lacks the capability for feature fine-tuning within the specific downstream tasks, limiting its adaptability and performance. To address this issue, we propose a Re-embedded Regional Transformer (R2T) for re-embedding the instance features online, which captures fine-grained local features and establishes connections across different regions. Unlike existing works that focus on pre-training powerful feature extractor or designing sophisticated instance aggregator, R2T is tailored to re-embed instance features online. It serves as a portable module that can seamlessly integrate into mainstream MIL models. Extensive experimental results on common computational pathology tasks validate that: 1) feature re-embedding improves the performance of MIL models based on ResNet-50 features to the level of foundation model features, and further enhances the performance of foundation model features; 2) the R2T can introduce more significant performance improvements to various MIL models; 3) R2T-MIL, as an R2T-enhanced AB-MIL, outperforms other latest methods by a large margin. The code is available at: https://github.com/DearCaat/RRT-MIL.
@@ 1 Introduction
In this paper, we propose Re-embedded Regional Transformer (R2T), which leverages the advantages of the native MSA while overcoming its limitations. Specifically, R2T applies the native MSA to each local region separately. Then, it uses a Cross-region MSA (CR-MSA) to fuse the information from different regions. Finally, a novel Embedded Position Encoding Generator (EPEG) is used to effectively encode the positional information of the patches. By incorporating with mainstream MIL models, the proposed R2T can ensure efficient computation while maintaining powerful representation capabilities to fine-tune the offline features according to the specific downstream tasks. The main contributions can be summarized as follows:
@@
For re-embedding instance features, we design a Re-embedded Regional Transformer (R2T) which can be seamlessly integrated into mainstream MIL models to further improve performance. By incorporating the R2T into AB-MIL, we present the R2T-MIL, which achieves state-of-the-art performance on various computational pathology benchmarks.
We introduce two novel components for the R2T: the Cross-region MSA and the Embedded Position Encoding Generator. The former enables effective information fusion across different regions. The latter combines the benefits of relative and convolutional position encodings to encode the positional information more effectively.
@@ 3.2 Re-embedded Regional Transformer
As illustrated in Figure 2, we propose a Re-embedded Regional Transformer (R2Transformer, R2T) to re-embed the input instance features as the new instance representations,
where is the mapping function of the R2Transformer here, is the processed input instance features, and is the new embedding of the -th instance. is the dimension of the embedding. The R2Transformer can be flexibly plugged into the MIL framework as a re-embedding module after feature input and before instance aggregation to reduce bias caused by the shift between offline feature learning and downstream tasks. The whole re-embedding process in the R2Transformer can be formulated as,
@@
| Methods | CAMELYON-16 | TCGA-BRCA | ||||||
| Accuracy | AUC | F1-score | Accuracy | AUC | F1-score | |||
| ResNet-50 ImageNet pretrained | AB-MIL [14] | 90.060.72 | 94.540.30 | 87.830.83 | 86.414.92 | 91.102.52 | 81.644.71 | |
| CLAM [22] | 90.140.85 | 94.700.76 | 88.100.63 | 85.172.70 | 91.671.78 | 80.373.04 | ||
| DSMIL [18] | 90.171.02 | 94.570.40 | 87.651.18 | 87.202.69 | 91.581.33 | 82.412.92 | ||
| TransMIL [26] | 89.222.32 | 93.512.13 | 85.104.33 | 84.682.67 | 90.801.91 | 79.862.63 | ||
| DTFD-MIL [45] | 90.220.36 | 95.150.14 | 87.620.59 | 85.921.76 | 91.431.64 | 81.092.05 | ||
| IBMIL [20] | 91.230.41 | 94.801.03 | 88.800.89 | 84.193.40 | 91.012.32 | 79.453.42 | ||
| MHIM-MIL [30] | 91.810.82 | 96.140.52 | 89.940.70 | 86.735.59 | 92.361.58 | 82.435.47 | ||
| R2T-MIL | 92.400.31 | 97.320.29 | 90.630.45 | 88.330.67 | 93.171.45 | 83.700.95 | ||
| PLIP WSI pretrained | AB-MIL [14] | 94.660.42 | 97.300.31 | 93.290.54 | 85.452.32 | 91.732.26 | 80.602.66 | |
| CLAM [22] | 93.730.54 | 97.170.50 | 91.600.60 | 86.701.35 | 92.162.02 | 81.911.78 | ||
| DSMIL [18] | 94.400.85 | 97.060.56 | 92.781.15 | 87.252.70 | 91.801.67 | 82.182.28 | ||
| TransMIL [26] | 94.400.43 | 97.880.21 | 92.810.43 | 85.833.44 | 92.172.20 | 81.123.25 | ||
| DTFD-MIL [45] | 94.570.31 | 97.290.23 | 93.120.40 | 86.422.67 | 92.162.42 | 81.772.73 | ||
| IBMIL [20] | 93.900.66 | 97.040.18 | 92.440.64 | 87.571.48 | 91.711.74 | 82.782.02 | ||
| MHIM-MIL [30] | 95.320.31 | 97.790.15 | 94.130.42 | 87.072.20 | 93.172.00 | 82.482.50 | ||
| R2T-MIL | 95.490.00 | 98.050.29 | 94.290.04 | 88.823.22 | 93.801.24 | 84.553.55 | ||
@@ 3.3 R2Transformer-based MIL
where denotes any MIL loss. R2Transformer-based MIL (R2T-MIL) adopts the instance aggregation method and the bag classifier of AB-MIL [14] by default.
@@ 4.2.1 Cancer Diagnosis, and Sub-typing
Table 1 presents the diagnosis and sub-typing performances of different MIL approaches on the C16 and BRCA datasets. The results demonstrate that our proposed R2T-MIL achieves the best performance under all metrics on all benchmarks. Specifically, R2T-MIL gets 0.59%, 1.18%, and 0.69% performance gains over the second-best methods in Accuracy, AUC, and F1-score respectively on the C16 dataset. On the BRCA dataset, the AUC improvement is 0.69%. R2T-MIL employs the same aggregation and classification methods as AB-MIL. However, R2T-MIL significantly outperforms AB-MIL. It increases the AUC by 2.78% and 1.77% on the C16 and BRCA datasets, respectively. The sub-typing results on NSCLC in Table 2 support a similar observation. We attribute these substantial performance improvements to the additional re-embedding step based on our proposed R2T, which surpasses the performance of the foundation model (+0.02% AUC on C16, +1.37% on BRCA, +0.72% on NSCLC). In addition, we find that R2T can further enhance the features of the foundation model, achieving considerable improvement. This validates the effectiveness of re-embedding.
@@
| ResNet-50 | AB-MIL | 90.32±1.39 | 95.29±1.14 | 89.83±1.53 |
| CLAM | 90.52±2.08 | 95.37±1.08 | 90.08±1.97 | |
| DSMIL | 90.43±2.52 | 95.60±0.81 | 90.03±2.61 | |
| TransMIL | 90.04±1.86 | 94.97±1.11 | 89.94±1.73 | |
| DTFD-MIL | 89.85±1.53 | 95.55±1.47 | 89.60±1.67 | |
| IBMIL | 90.04±1.48 | 95.57±1.13 | 89.73±1.64 | |
| MHIM-MIL | 91.27±2.35 | 96.02±1.35 | 90.85±2.53 | |
| R2T-MIL | 91.75±2.38 | 96.40±1.13 | 91.26±2.60 | |
| PLIP | AB-MIL | 90.99±2.43 | 95.68±1.98 | 90.52±2.45 |
| CLAM | 90.80±2.35 | 95.46±1.72 | 90.38±2.46 | |
| DSMIL | 90.60±2.37 | 95.78±1.81 | 90.24±2.51 | |
| TransMIL | 89.09±2.00 | 95.30±1.95 | 88.83±2.16 | |
| DTFD-MIL | 90.42±2.98 | 95.83±1.75 | 89.91±3.01 | |
| IBMIL | 91.18±3.27 | 95.62±2.09 | 90.94±3.20 | |
| MHIM-MIL | 91.74±1.88 | 96.21±1.26 | 91.20±1.89 | |
| R2T-MIL | 92.13±2.55 | 96.40±1.45 | 91.83±2.50 | |
@@
| ResNet-50 | AB-MIL | 57.50±3.94 | 58.78±4.90 | 56.51±7.14 |
| CLAM | 57.57±3.73 | 59.60±3.93 | 56.65±6.90 | |
| DSMIL | 57.42±2.25 | 59.31±4.75 | 55.03±6.61 | |
| TransMIL | 58.90±4.70 | 64.11±1.99 | 56.39±2.94 | |
| DTFD-MIL | 56.98±3.24 | 59.48±2.61 | 55.16±4.33 | |
| IBMIL | 58.41±2.90 | 58.58±4.67 | 59.18±3.29 | |
| MHIM-MIL | 58.36±3.26 | 60.32±4.41 | 56.08±6.33 | |
| R2T-MIL | 61.13±2.36 | 67.19±4.02 | 60.95±4.41 | |
| PLIP | AB-MIL | 59.18±2.48 | 62.09±4.38 | 57.12±2.39 |
| CLAM | 61.58±2.89 | 64.05±4.70 | 58.00±3.34 | |
| DSMIL | 58.96±1.80 | 63.82±5.56 | 56.12±2.21 | |
| TransMIL | 56.20±3.26 | 63.55±2.94 | 58.84±3.28 | |
| DTFD-MIL | 59.67±4.71 | 61.78±2.33 | 57.75±3.52 | |
| IBMIL | 56.32±2.69 | 58.86±3.40 | 57.33±3.28 | |
| MHIM-MIL | 60.92±3.38 | 62.94±4.58 | 55.95±2.54 | |
| R2T-MIL | 63.98±2.26 | 65.94±1.34 | 60.42±2.15 | |
@@ 4.2.2 Survival Prediction
Table 3 shows the experimental results on three survival prediction datasets. It is worth noting that our proposed R2T-MIL model demonstrates outstanding performance, attaining a C-index of 61.13% on the BLCA, 67.19% on the LUAD, and 60.95% on the LUSC. It outperforms the compared methods by a significant margin, with improvements of 2.23%, 3.08%, and 1.77% over the second-best methods, respectively. Furthermore, our proposed feature re-embedding strategy can yield substantial improvements even when working with high-quality features extracted by the foundation model. Particularly, compared to AB-MIL, our feature re-embedding strategy brings performance improvements of 4.8%, 3.85%, and 3.3% for the three datasets, respectively. These results highlight the consistent and reliable performance of our proposed strategy and method, indicating its efficacy in predicting survival outcomes.
@@ 4.3.1 Re-Embedding Matters in Pathology
Foundation Model Features vs. Re-embedding Features: Table 4 and Figure 4 compare the performance of the features extracted by the foundation model PLIP [13] and various re-embedding features. The PLIP is based on the multi-modal foundation model CLIP [24] and uses up to 200K slide-text pairs for pre-training. Although this high-cost pre-training brings some improvement on different tasks, it still has bottlenecks. We attribute this to the dilemma of the traditional paradigm that even the best offline pre-training features cannot address the issue of insufficient feature fine-tuning for downstream tasks. In contrast, re-embedding modules that can be end-to-end trained with MIL models provide supervised feature fine-tuning, which enables full exploitation of the knowledge beneficial to the final task. Hence, we can see from Table 4 that any re-embedding structure achieves a considerable improvement on all tasks. Some tailored structures, such as our proposed R2T, can have a significant performance advantage in most of the tasks. Moreover, Figure 4 shows that this improvement is not limited to the classical AB-MIL, but can widely benefit different MIL models. Table 1 demonstrates that the re-embedding approach is still effective on foundation model features. Therefore, compared to foundation model features, re-embedding features are a cheaper, more versatile alternative and effective booster.
@@
| +R2T (local) | 97.32 | 96.40 | 67.19 | 6.5s |
Different Re-embedding Discussion: The bottom part of Table 4 presents the performance of AB-MIL under different feature settings. We employ four methods, including TranMIL, N-MSA, the local version of N-MSA, and our proposed R2T to re-embed the features. From observations, all four re-embedded AB-MILs perform better than the original one. The performance improvement by TransMIL, N-MSA, and R2T are 1.28%, 1.68%, and 2.80%, respectively, on C16, while these numbers on LUAD are 4.46%, 4.14%, and 8.41%, respectively. This phenomenon validates the importance of re-embedding in MIL-based computational pathology. Among the four employed re-embedding approaches, R2T boosts the AB-MIL with the most considerable improvements while incurring the lowest computational cost. Specifically, R2T achieves more performance gains (+0.85% on C16 and +2.63% on LUAD) while only requiring a 1/5 inference time compared with the second-best approaches.
The Applicability of R2T in MIL Frameworks: We incorporate the R2T into different MIL frameworks as a re-embedding module for studying its applicability. The performance improvements achieved by this re-embedding module in different frameworks are shown in Figure 4. The results reveal that the R2T is capable of improving all MIL frameworks on all tasks. Moreover, the improvement brought by R2T surpasses the foundation model features on two tasks except for diagnosis, and reaches a similar level on C16. This clearly verifies the good applicability of R2T.